OpenStudy (anonymous):

medal & fan Which of the following best defines 4 to the power of 2 over 3?

OpenStudy (anonymous):

Square root of 16 Cube root of 4 Square root of 4 Cube root of 16

OpenStudy (anonymous):

\[\large (4)^{\frac{2}{3}}\]

OpenStudy (anonymous):

Use this rule: \[\large \color{green}{(x)^{\frac{m}{n}} = \sqrt[n]{(x)^m}}\]

OpenStudy (anonymous):

The moment you succeed to use this rule, you will get your answer.. :)

OpenStudy (anonymous):

So it would be the square root of 4?

OpenStudy (anonymous):

cube root of 16 i think

OpenStudy (anonymous):

Really?? I said use that rule and write here what you got, Simplify we can do later..

OpenStudy (anonymous):

@mason437 What you think does not matter here, what the asker thinks, is important.. :P

OpenStudy (anonymous):



OpenStudy (anonymous):